19. Change 6 quarts to an equivalent role day in pints ? nour DIMENSIONAL ANALYSIS REQUIRED
Answer #1

Given : 6 quarts/ day

Quart and Pint are the unit of liquid capacity.

Pint is a Solution :

Time 1 day = 24 hours ( \therefore 6 quarts in 24 hours )

1 quart = 2 pints

To find: quats/day to pints / hour

Dimensional analysis:

unit you have * conversion factor = unit you want

6 quarts / 24 hours * 2 pints = X pints / hour

\therefore 6 * 2 /24 = 12 / 24 = 1/2 = 0.5 pints / hour

Ans : 6 quarts day = 0.5 pints / hour